A Bibliography of Supercomputing ’98

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 2.02

Abstract

This bibliography records articles presented at the Supercomputing ’98 conference.

Title word cross-reference

2000 [FHJ^{+}98].

3D [LYCD98].

Accessed [CDI98]. Accurate [TS98]. ACM/IEEE [ACM98]. Adaptive [Mar98]. Adaptively [SKK98]. AGCM [MDFS98]. Agent [HIIS98]. Algebra [MvdG98, WD98]. Algorithm [PHS98]. Algorithms [BF98, KK98, PPR98, SS98a]. Analysis [BGC98, GC98]. Analyzing [SGS98]. Application [LYCD98]. Applications [HH98]. Architecture [BTE98, BGC98, Gar98]. Architectures [QB98]. Array [PPR98]. ATLAS [WD98]. Authoring [AFFH98]. Automatic [FHJ^{+}98]. Automatically [WD98]. Available [CDI98].

Balancing [KC98]. Barrier [LJ98]. Based [SGS98, SSCM98, TS98]. Benchmark [BTE98]. Beowulf [SSS98]. Beowulf-class [SSS98]. Between [PHS98, PPR98]. Block [PPR98]. Block-Cyclic [PPR98]. Bounds [SGS98].

C3I [BTE98]. Cache [QB98]. Cache-Coherent [QB98]. Case [TGL98]. CCF [Sun98]. Center [ACM98]. Challenge [Mar98]. class [SSS98]. Cluster [QB98]. Code [TS98]. Codes [AJMCY98]. Coherence [LJ98]. Coherent [QB98]. Collaborative [PTC98, Sun98]. Collective [SCHC98]. Combustion [WEW98].

Communication [ABD^{+}98, BF98]. Comparison [FHJ^{+}98]. Compilation [AJMCY98]. Computation [Mar98]. Computational [SSCM98]. Computations [KC98]. Computers [SS98a]. Computing [ACM98, AAN^{+}98, AFFH98, BBD^{+}98, FK98, MDFS98, MCF^{+}98, Sun98]. Condominium [AAN^{+}98]. Conference [ACM98]. Configuration [DNH98]. Conformational [LYCD98]. Constraint [KK98]. Convention [ACM98]. core [DNH98]. Corona [Mar98]. County [ACM98]. Cycle [RH98]. Cyclic [PPR98].

Data [CDI98, GC98]. Datatypes [TGL98]. Delivering [HH98]. Derived [TGL98]. Design [Gar98]. Distributed [AFFH98, FK98, PTC98, SSS98, SS98a]. DSMs [LJ98]. Dynamic [CDI98, KC98, NB98, SKK98, SS98b].

Efficiency [TCL98]. Efficient [Gar98, HTI98, MCF^{+}98, PPR98, SS98a]. Eigensolver [SSH98]. Ejection [Mar98]. Electronic [WCS98]. Elimination [LD98]. Enabled [FK98]. Engines [WEW98]. Environment [AFFH98, SSCM98]. Environments [MDFS98, MCF^{+}98, MvdG98]. Error [SGS98]. Evaluation [BTE98, TS98]. Execution [MCF^{+}98]. Exploring [ACR^{+}98].

File [Gar98]. Fine [RH98]. Fine-Grain [RH98]. Florida [ACM98]. Fortran [AJMCY98]. Framework [BBD^{+}98, KC98, PTC98]. Frameworks [Sun98]. Frequently [CDI98].

Gang [HTI98]. Gaussian [LD98]. Grain [RH98]. Grand [Mar98]. Grand-Challenge [Mar98]. Graph [KK98]. Grid [FK98]. Grid-Enabled [FK98]. Grids [PHS98]. Guided [WJK98].

Hardware [Ber98]. HARP [SS98b]. Heterogeneous [FK98, HIIS98]. Hierarchical [KC98]. High [ACM98, AFFH98, BBD^{+}98, GC98, MDFS98, SSH98, WEW98, AJMCY98]. High-Level [AFFH98]. High-Performance [BBD^{+}98, WEW98]. Highly [CDI98, Gar98, HTI98].

I/O [SCHC98, TGL98]. IBM [Ber98].

Image [WJK98]. Image-Guided [WJK98]. Implementation [BGC98, DNH98, HTI98]. Improve [TGL98]. Infrastructure [MCF^{+}98]. Initial [BTE98]. Interaction [DNH98, HHC^{+}98]. Interface [BGC98]. Internal [WEW98]. Interpolation [PHS98]. Irregular [NB98]. Issues [ACR^{+}98].

Job [MCF^{+}98].

Lanczos [WCS98]. Large [LYCD98]. Level [AFFH98]. Libraries [SCHC98]. Lightweight [SCHC98]. like [MvdG98]. Linear [MvdG98, WD98]. Load [KC98]. Load-Balancing [KC98]. Loops [GN98].

Making [LD98]. Mass [Mar98]. Massively [DNH98]. MATLAB [MvdG98]. MATLAB-like [MvdG98]. Matrix [TCL98]. Measurement [Ber98]. Memory [SS98a, TCL98]. Meshes [SKK98]. Message [FK98]. Method [WCS98]. Microscopy [PTC98]. Middleware [HIIS98, TS98]. Mining [GC98]. Modelling [TS98]. Molecular [TS98]. Monitor [Ber98]. MPI [FK98, HH98, TGL98]. MPI-StarT [HH98]. MTA [SCB^{+}98]. Muli [KK98]. Muli-Constraint [KK98]. Multi [BF98, SCB^{+}98]. Multi-processor [SCB^{+}98]. Multi-Tier [BF98]. Multidimensional [GC98]. Multilevel [KK98]. Multiple [PHS98]. Multiplication [TCL98]. Multipole [SGS98].

Multipole-Based [SGS98]. Multireference [DNH98]. Multithreaded [BTE98, KC98].

Network [HH98]. Networking [ACM98]. Networks [LHZ98, RH98]. Notebook [SSCM98]. Notebook-Based [SSCM98]. November [ACM98]. Numerical [HH98].

O [SCHC98, TGL98]. Object [BBD^{+}98]. Object-Oriented [BBD^{+}98]. OpenMP [LHZ98]. Optimizing [QB98]. Orange [ACM98]. Oriented [BBD^{+}98]. Origin [FHJ^{+}98]. Orlando [ACM98]. Out-of-core [DNH98]. Overhead [LJ98]. Overlap [BF98]. OVERTURE [BBD^{+}98].

Parallel [BF98, BTE98, DNH98, Gar98, HHC^{+}98, Mar98, MCF^{+}98, MvdG98, PHS98, SSH98, SS98b, SCHC98]. Parallelism [NB98]. Parallelization [FHJ^{+}98, GN98, LYCD98]. Parallelizing [AJMCY98]. Partitioner [SS98b]. Partitioning [KK98]. Passing [FK98]. Path [Mar98]. PDS [SCHC98]. PDS/PIO [SCHC98]. Compilers [FHJ^{+}98]. IEEE [ACM98]. PIO [SCHC98]. Penn [AAN^{+}98]. Performance [ACM98, AJMCY98, AFFH98, BBD^{+}98, Ber98, GC98, HH98, MDFS98, SSH98, SCB^{+}98, TS98, TGL98, WEW98]. PFS [Gar98]. Pivoting [LD98]. Prediction [LYCD98, Mar98, TS98]. Problem [Mar98]. Proceedings [ACM98]. Processor [PPR98, SCB^{+}98]. Program [DNH98]. Programming [AFFH98, BTE98]. Protein [LYCD98]. Prototype [SSCM98]. Pthreads [NB98].

Quantitative [ABD^{+}98].

Real [WJK98]. Real-Time [WJK98]. Redistribution [PPR98]. Reducing [LJ98]. Refined [SKK98]. Rendezvous [PHS98].

Repartitioning [SKK98]. Run [GN98]. Run-Time [GN98]. Runtime [HHC^{+}98].

S [SS98b]. S-HARP [SS98b]. SC98

[ACM98, ACM98]. Scalable

[ACR^{+}98, CDI98, LD98]. Scale [LYCD98]. Scaling [SSS98]. Scheduling [AAN^{+}98, HTI98]. Scientific [AJMCY98, BBD^{+}98, Ber98, HIIS98]. Segmentation [WJK98]. Selection [SS98a]. Server [MvdG98]. Serving [CDI98]. Sets [PPR98]. SGI [FHJ^{+}98]. Simulation [TS98, WEW98, WCS98]. Simulations [HHC^{+}98, HIIS98]. Sites [CDI98]. SMP [ACR^{+}98]. Software [LJ98, QB98, WD98]. Solar [Mar98]. Solar-Corona [Mar98]. Space [ABD^{+}98]. Sparse [LD98]. Spectral [SS98b]. Speculative [GN98]. StarT [ACR^{+}98, HH98]. StarT-Voyager [ACR^{+}98]. State [AAN^{+}98]. Static [LD98]. Stealing [RH98]. Strassen [TCL98]. Structure [LYCD98, WCS98]. Study [ABD^{+}98]. Suite [BTE98]. Supporting [HHC^{+}98]. Surgery [WJK98]. Synchronization [LJ98]. System [BTE98, CDI98, Gar98, AAN^{+}98]. Systems [FK98, SSS98].

Techniques [AJMCY98, GN98]. Tera [BTE98, SCB^{+}98]. Terascale [MCF^{+}98]. TFLOPS [Gar98]. Tier [BF98]. Time [GN98, WJK98]. Tool [HHC^{+}98]. Toolkit [AFFH98]. Tools [FHJ^{+}98, SSCM98]. Tools/Compilers [FHJ^{+}98]. Treecodes [SGS98]. Tuned [WD98]. Tuning [TCL98].

UCLA [MDFS98]. USA [ACM98]. User [ABD^{+}98]. User-Space [ABD^{+}98]. Using [Ber98, BTE98, TGL98].

Virtual [BGC98]. Visual [AFFH98].

Voyager [ACR^{+}98].

Web [CDI98]. WebFlow [AFFH98]. Workload [Ber98]. Workstations [LHZ98, RH98].

References

Agnihotri:1998:PSC

[AAN^{+}98] Pawan Agnihotri, Vijay K. Agarwala, Jeffrey J. Nucciarone, Kevin M. Morooney, and Chita Das. The Penn State Computing Condominium Scheduling System. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Agarwala833/
index.htm.

Araki:1998:USC

[ABD^{+}98] Soichiro Araki, Angelos Bilas, Cezary Dubnicki, Jan Edler, Koichi Konishi, and James Philbin. User-space communication: A quantitative study. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Bilas820/index.
htm.

ACM:1998:SHP

[ACM98] ACM, editor. SC’98: High Performance Networking and Computing: Proceedings of the 1998 ACM/IEEE SC98 Conference: Orange County Convention Center, Orlando, Florida, USA, November 7–13, 1998. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN ????

LCCN ???? URL http://www.
supercomp.org/sc98/papers/.

Ang:1998:SVE

[ACR^{+}98] Boon S. Ang, Derek Chiou, Daniel L. Rosenband, Mike Ehrlich, Larry Rudolph, and Arvind. StarT-Voyager: Exploring scalable SMP issues. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Ang1085/index.
htm.

Akarsu:1998:WHL

[AFFH98] Erol Akarsu, Geoffrey C. Fox, Wojtek Furmanski, and Tomasz Haupt. WebFlow — high-level programming environment and visual authoring toolkit for high performance distributed computing. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Akarsu809/Index.
htm.

Adve:1998:HPF

[AJMCY98] Vikram S. Adve, Guohua Jin, John Mellor-Crummey, and Qing Yi. High Performance Fortran compilation techniques for parallelizing scientific codes. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Adve1294/INDEX.
HTM.

Bassetti:1998:OOO

[BBD^{+}98] Federico Bassetti, David Brown, Kei Davis, William Henshaw, and Dan Quinlan. OVERTURE: An object-oriented framework for high-performance scientific computing. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Bassetti949/
INDEX.HTM.

Bergeron:1998:MSW

[Ber98]  Robert J. Bergeron.  Mea-

surement of a scientific workload using the IBM hardware performance monitor. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Bergeron1289/
index.htm.

Baden:1998:COM

[BF98] Scott B. Baden and Stephen J.

Fink. Communication overlap in multi-tier parallel algorithms. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Baden708/INDEX.
HTM.

Buonadonna:1998:IAV

[BGC98] Philip Buonadonna, Andrew Geweke, and David Culler. Implementation and analysis of the virtual interface architecture. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Buonadonna893/
INDEX.HTM.

Brunett:1998:IET

[BTE98] Sharon M. Brunett, John Thornley, and Marrq Ellenbecker. An initial evaluation of the Tera multithreaded architecture and programming system using the C3I parallel benchmark suite. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Brunett1063/
Index.htm.

Challenger:1998:SHA

[CDI98] Jim R. Challenger, Paul Dantzig, and Arun Iyengar. A scalable and highly available system for serving dynamic data at frequently accessed Web sites. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Challenger602/
index.htm.

Dachsel:1998:CIM

[DNH98] Holger Dachsel, Jarek Nieplocha, and Robert Harrison. An outof-core implementation of the massively parallel multireference configuration interaction program. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Dachsel897/
INDEX.HTM.

Frumkin:1998:CAP

[FHJ^{+}98] Michael Frumkin, Michelle Hribar, Haoqiang Jin, Abdul Waheed, and Jerry Yan. A comparison of automatic parallelization tools/compilers on the SGI Origin 2000. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Hribar1140/
INDEX.HTM.

Foster:1998:GEM

[FK98] Ian Foster and Nicholas T. Karonis. A grid-enabled MPI: Message passing in heterogeneous distributed computing systems. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Foster1125/
index.htm.

Garg:1998:TPA

[Gar98] Sharad Garg. TFLOPS PFS:

Architecture and design of a highly efficient parallel file system. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Garg891/index.
htm.

Goil:1998:HPM

[GC98] Sanjay Goil and Alok Choudhary. High performance multidimensional analysis and data mining. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.

                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Goil660/index.
                                        htm.

Gupta:1998:TSR

[GN98] Manish Gupta and Rahul Nim. Techniques for speculative runtime parallelization of loops. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Gupta535/index.
                                        htm.

Husbands:1998:MSD

[HH98] Parry J. Husbands and James C. Hoe. MPI-StarT: Delivering network performance to numerical applications. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Husbands924/
                                        INDEX.HTM.

Harrop:1998:SRT

[HHC^{+}98] Christopher W. Harrop, Steven T. Hackstadt, Janice E. Cuny, Allen D. Malony, and Laura S. Magde. Supporting runtime tool interaction for parallel simulations. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Harrop888/index.
                                        htm.

Ho:1998:AMH

[HIIS98] Shirun Ho, Satoshi Itoh, Sigeo Ihara, and Richard D. Schlichting. Agent middleware for heterogeneous scientific simulations. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Ho693/index.
htm.

Hori:1998:HEG

[HTI98] Atshushi Hori, Hiroshi Tezuka, and Yutaka Ishikawa. Highly efficient gang scheduling implementation. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Hori698/INDEX.
HTM.

Karamcheti:1998:HLB

[KC98] Vijay Karamcheti and Andrew A. Chien. A hierarchical load-balancing framework for dynamic multithreaded computations. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Karamcheti553/
index.htm.

Karypis:1998:MAM

[KK98] George Karypis and Vipin Kumar. Multilevel algorithms for muli-constraint graph partitioning. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Karypis908/
index.htm.

Li:1998:MSG

[LD98] Xiaoye Sherry Li and James W. Demmel. Making sparse Gaussian elimination scalable by static pivoting. In ACM [ACM98], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.supercomp.org/sc98/
                                        TechPapers/sc98_FullAbstracts/
                                        Li844/index.htm.

Lu:1998:ONW

[LHZ98] Honghui Lu, Y. Charlie Hu, and Willy Zwaenepoel. OpenMP on networks of workstations. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Lu1105/index.
                                        htm.

Lee:1998:RCO

[LJ98] Jae Bum Lee and Chu Shik Jhon. Reducing coherence overhead of barrier synchronization in software DSMs. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Lee583/INDEX.
                                        HTM.

LoCascio:1998:LSP

[LYCD98] Philip F. LoCascio, Kaizhi Yue, Peter Cummings, and Ken Dill. The large scale parallelization of a conformational 3D protein structure prediction application. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_

FullAbstracts/LoCascio1021/
index.htm.

Marshall:1998:APC

[Mar98] Hal G. Marshall. Adaptive parallel computation of a grandchallenge problem: Prediction of the path of a solar-corona mass ejection. In ACM [ACM98], page ?? ISBN ???? LCCN

????  URL http://www.
supercomp.org/sc98/papers/.

Moreira:1998:IEP

[MCF^{+}98] Jose E. Moreira, Waiman Chan, Liana L. Fong, Hubertus Franke, and Morris A. Jette. An infrastructure for efficient parallel job execution in terascale computing environments. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Moreira808/
index.htm.

Mechoso:1998:UAH

[MDFS98] Carlos R. Mechoso, L. A. Drummond, J. D. Farrara, and J. A. Spahr. The UCLA AGCM in high performance computing environments. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Mechoso916/
INDEX.HTM.

Morrow:1998:PLA

[MvdG98] Greg Morrow and Robert van de Geijn. A parallel linear algebra server for MATLAB-like environments. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Morrow779/index.
                                        htm.

Narlikar:1998:PDI

[NB98] Girija J. Narlikar and Guy E. Blelloch. Pthreads for dynamic and irregular parallelism. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Narlikar821/
                                        index.htm.

Plimpton:1998:PRA

[PHS98] Steve Plimpton, Bruce Hendrickson, and James Stewart. A parallel rendezvous algorithm for interpolation between multiple grids. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Plimpton644/
                                        index.htm.

Park:1998:EAB

[PPR98] Neungsoo Park, Viktor K. Prasanna, and Cauligi Raghavendra. Efficient algorithms for block-cyclic array redistribution between processor sets. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Park1090/index.
                                        htm.

Parvin:1998:CFD

[PTC98] B. Parvin, J. Taylor, and G. Cong. A collaborative framework for distributed microscopy. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Parvin1020/
index.htm.

Qin:1998:OSC

[QB98] Xiaohan Qin and Jean-Loup Baer. Optimizing software cachecoherent cluster architectures. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Qin753/index.
htm.

Ryu:1998:FGC

[RH98] Kyung D. Ryu and Jeffrey K.

Hollingsworth. Fine-grain cycle stealing for networks of workstations. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Ryu647/index.
htm.

Snavely:1998:MPP

[SCB^{+}98] Allan E. Snavely, Larry Carter, Jay Boisseau, Amit Majumdar, Kang Su Gatlin, Nick Mitchell, John Feo, and Brian Koblenz. Multi-processor performance on the Tera MTA. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Snavely973/
INDEX.HTM.

Sturtevant:1998:PPL

[SCHC98] Judy E. Sturtevant, Mark Christon, Philip D. Heermann, and Pang-Chieh Chen. PDS/PIO: Lightweight libraries for collective parallel I/O. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Sturtevant894/
                                        index.htm.

Sarin:1998:AEB

[SGS98] Vivek Sarin, Ananth Grama, and Ahmed Sameh. Analyzing the error bounds of multipole-based treecodes. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Sarin1042/index.
                                        htm.

Schloegel:1998:DRA

[SKK98] Kirk A. Schloegel, George Karypis, and Vipin Kumar. Dynamic repartitioning of adaptively refined meshes. In ACM [ACM98], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.supercomp.org/sc98/
                                        TechPapers/sc98_FullAbstracts/
                                        Schloegel995/INDEX.HTM.

Saukas:1998:ESA

[SS98a] Einar L. G. Saukas and S. W. Song. Efficient selection algorithms on distributed memory computers. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_

FullAbstracts/Saukas485/INDEX.
HTM.

Sohn:1998:HPD

[SS98b] Andrew Sohn and Horst Simon. S-HARP: A parallel dynamic spectral partitioner. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Sohn525/index.
htm.

Skidmore:1998:PNB

[SSCM98] Jenifer Skidmore, Matthew J.

Sottile, Janice E. Cuny, and Allen D. Malony. A prototype notebook-based environment for computational tools. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Skidmore859/
INDEX.HTM.

Sears:1998:HPP

[SSH98] Mark P. Sears, Ken Stanley, and Greg Henry. A high performance parallel eigensolver. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Sears974/index.
htm.

Salmon:1998:SBC

[SSS98] John K. Salmon, Christopher Stein, and Thomas Sterling. Scaling of Beowulf-class distributed systems. In ACM [ACM98], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.supercomp.org/sc98/
                                        TechPapers/sc98_FullAbstracts/
                                        Salmon793/index.htm.

Sunderam:1998:CCC

[Sun98] Vaidy Sunderam. CCF: Collaborative computing frameworks. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Sunderam1290/
                                        index.htm.

Thottethodi:1998:TSM

[TCL98] Mithuna S. Thottethodi, Siddhartha Chatterjee, and Alvin R. Lebeck. Tuning Strassen’s matrix multiplication for memory efficiency. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Thottethodi872/
                                        index.htm.

Thakur:1998:CUM

[TGL98] Rajeev S. Thakur, William Gropp, and Ewing Lusk. A case for using MPI’s derived datatypes to improve I/O performance. In ACM [ACM98], page ?? ISBN ???? LCCN ????

                                        URL http://www.supercomp.
                                        org/sc98/TechPapers/sc98_
                                        FullAbstracts/Thakur447/index.
                                        htm.

Taufer:1998:APE

[TS98] Michela Taufer and Thomas Stricker. Accurate performance evaluation, modelling and prediction of a molecular simulation code based on middleware. In ACM [ACM98], page ?? ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Taufer597/INDEX.
HTM.

Wu:1998:NLM

[WCS98] Kesheng Wu, A. Canning, and H. D. Simon. A new Lanczos method for electronic structure simulation. In ACM [ACM98], pages 35–40. ISBN ???? LCCN ???? URL http:

//www.supercomp.org/sc98/
TechPapers/sc98_FullAbstracts/
Wu1023/index.htm.

Whaley:1998:ATL

[WD98] R. Clint Whaley and Jack J.

Dongarra. Automatically Tuned Linear Algebra Software (AT-LAS). In ACM [ACM98], page ???? ISBN ???? LCCN ???? URL http:/

/www.netlib.org/utk/people/
JackDongarra/PAPERS/atlas-
sc98.ps;   http://www.supercomp.
org/sc98/TechPapers/sc98_
FullAbstracts/Whaley814/INDEX.

HTM. Best Paper Award for Systems.

Werner:1998:HPS

[WEW98] Andreas Werner, Harald Echtle, and Monika Wierse. Highperformance simulation of internal combustion engines. In ACM [ACM98], page ?? ISBN ???? LCCN ???? URL http:

//www.supercomp.org/sc98/
TechPapers/sc98_FullAbstracts/
Werner772/index.htm.

Warfield:1998:RTI

[WJK98] Simon K. Warfield, Ferenc A. Jolesz, and Ron Kikinis. Realtime image segmentation for image-guided surgery. In ACM [ACM98], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.supercomp.org/sc98/
                                        TechPapers/sc98_FullAbstracts/
                                        Warfield1114/index.htm.